A) According to a survey it takes .7 second to download the home page website. Suppose that the download was normally distributed with a standard deviation of .2.

If you select a random sample of 23 download times, the probability is 85% that the sample mean is between what two values symmetrically distributed around the population mean?

Lower Bound:

Upper Bound:
Answer #1

Here confidence level = 85%

So, alpha = 0.15

Now, z_{alpha/2} = 1.44

Lower Bound

= i-Zo/2- 0.7- 1.44 Vn V 23 = 0.6399 0.6399

Upper Bound

0.2 = x + a /2-= 0.7 + 1.44 * 0.7601 Vn V23

** If the answer does not match please comment